OpenStudy (anonymous):

MEDAL FOR HELP Solve for x: -2x - 6 = -12

OpenStudy (anonymous):

x = -3, x = 4.5 x = 3, x = 9 x = -3, x = 9 No solutions

jigglypuff314 (jigglypuff314):

to find x for -2x - 6 = -12 first add 6 to both sides +6 +6 -2x = -6 then divide both sides by -2 /-2 /-2 x = -6/-2 = 6/2 so x = ^

OpenStudy (anonymous):

jigglypuff i guess you could handle it ..

OpenStudy (anonymous):

@jigglypuff314 I messed up this was the quesiton Solve for x: -2|x - 3| = -12 x = -3, x = 4.5 x = 3, x = 9 x = -3, x = 9 No solutions

jigglypuff314 (jigglypuff314):

to solve for x for -2|x - 3| = -12 first isolate the absolute values /-2 /-2 so divide both sides by -2 you get |x - 3| = 6 which you can separate into x - 3 = 6 and -(x - 3) = 6 then solve for x +3 +3 - x + 3 = 6 x = 6 + 3 x = 3 - 6 then simplify

OpenStudy (anonymous):

So the third choice? @jigglypuff314

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

You are awesome and do you think you got time for 1 more?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Solve: 4x - 2(x + 1) = 3x + 10 -12 12 -8 8

jigglypuff314 (jigglypuff314):

to solve for x for 4x - 2(x + 1) = 3x + 10 use distributive property 4x + (-2)(x) + (-2)(1) = 3x + 10 simplify 4x - 2x - 2 = 3x + 10 2x - 2 = 3x + 10 then minus 10 from both sides -10 -10 2x - 12 = 3x then minus 2x from both sides -2x -2x -12 = x

OpenStudy (anonymous):

Thank you so much I can't thank you enough @jigglypuff314

jigglypuff314 (jigglypuff314):

your welcome :)

OpenStudy (anonymous):

Alright well I got to run thanks again @jigglypuff314 Also if you got skype maybe add me? Mines Bluetydestudios